Can you help me find the slope and y-intercept of X+Y=8 and the other problem is 2x+y=6 ?
Answer by stanbon(75887)

1) x+y=8
Put this in the form y=mx+b
Then m will be the slope, and
b will be the y-intercept.
So, y=-x+8
The slope = m =1, and
the y-intercept is "8".
2) 2x+y=6
So, y=-2x+6
slope = m = -2
y-intercept = 6
Cheers,
Stan H.